# SSL4EO-S12 v1.1: A Multimodal, Multiseasonal Dataset for Pretraining, Updated

### v1

@@ Abstract
This technical report presents SSL4EO-S12 v1.1, a multimodal, multitemporal Earth Observation dataset designed for pretraining large-scale foundation models. Building on the success of SSL4EO-S12 v1.0, the new version addresses the previous challenges of data misalignment and a limited data structure for low-barrier, analysis-ready EO processing. SSL4EO-S12 v1.1 covers the world’s 10,000 largest cities and its surroundings within a 50 km radius across four seasons, resulting in a diverse collection of nearly one million patches. SSL4EO-S12 v1.1 packages the data in Zarr file format for cloud-efficient loading and representation of meta-information such as including cloud masks and geolocation. Released under the CC-BY-4.0 license, SSL4EO-S12 v1.1 facilitates open research and provides a robust foundation for future advancements in self-supervised learning and geospatial analysis. The dataset is available online through https://datapub.fz-juelich.de/ssl4eo-s12, and we provided additional resources at https://github.com/DLR-MF-DAS/SSL4EO-S12-v1.1.
footnotetext: Co-funded by the European Union (Horizon Europe, Embed2Scale, 101131841).
@@ 1 Introduction
Figure 1: Example patches from SSL4EO-S12 v1.1 showing four columns with timestamps of Sentinel-2 (S2) L2A (left) and Sentinel-1 (S1) GRD (right) products. S1 GRD is visualized using a VH-$V$-VH pseudo coloring. $V$ is scaled from -30 – 5 while VH is scaled from -40 – 0 to account for their value ranges.
@@
Despite its success, the original SSL4EO-S12 dataset presented two main challenges: (1) misalignments between Sentinel-1 (S1) and Sentinel-2 (S2) modalities and (2) limited availability of analytics-ready data (ARD). To address these issues, we present SSL4EO-S12 v1.1. We ensure proper alignment by downloading larger areas and reprojecting S1 GRD data to match S2 L2A coordinates. Furthermore, we introduce stricter filtering and provide the dataset in an ARD format suitable for direct pretraining of EO FMs. Figure 1 illustrates samples from SSL4EO-S12 v.1.1.
This technical report is organized as follows: We first detail the data processing steps necessary to construct SSL4EO-S12 v1.1, including alignment, reprojection, and filtering. We then present a comprehensive overview of the final dataset and discuss its applicability for EO FM training pipelines.
@@ 2 Related Work
SSL4EO-S12 v1.0 addressed many of these limitations by including multiple modalities (e.g., S1 and S2) and covering multiple seasons, thereby supporting a broader range of downstream tasks [10]. Several state-of-the-art models have leveraged SSL4EO-S12 v1.0 to achieve significant improvements in accuracy and generalizability [10, 9, 4]. For instance, contrastive learning approaches and transformer-based architectures have demonstrated notable performance gains when pre-trained on SSL4EO-S12 v1.0.
@@ 3 Data Processing
All data for SSL4EO-S12 v1.1 were acquired from Google Earth Engine [5]. In v1.0, all samples were projected onto EPSG:4326, which can distort distances, especially near the poles. We retain the native UTM zones in v1.1 to preserve spatial accuracy and minimize unnecessary reprojections. For each location in v1, four timestamps from different seasons were selected from 2019 to 2021 with cloud coverage below 10% [10]. We selected the timestamp closest to the S2 timestamps in v1. To ensure correct spatial alignment, we first downloaded a larger area of interest for both S2 and S1 to buffer spatial mismatch on reprojection of these modalities when paired with other datasets. Next, we checked that all S2 timestamps for a location shared the same UTM zone. If timestamps spanned multiple zones—common at UTM boundaries–we picked the majority CRS for that location, cropped a reference timestamp to a common size of 264x264 pixels, and reprojected the remaining timestamps accordingly. Subsequently, we projected S1 patches onto the curated 264x264-pixel S2 patches. Following v1 [10], we fixed the patch sizes to 264 pixels as it is divisible by six, accommodating S2 pixel resolutions of 10 m, 20 m, and 60 m. All bands were upsampled to 10 m using nearest interpolation.
@@
Next, we dropped any location with more than 1% of missing values (NaN) for any timestamp. The remaining NaNs were filled by nearest neighbor interpolation. We split the dataset into training and validation sets, following a 99%-to-1% partition defined by the TerraMesh datasetTerraMesh is a larger pretraining dataset to be released in Q2 2025. to avoid spatial overlap that may leak information while training models. We add a numerical offset of +1000 to the data to facilitate the latest processing standard defined by ESA [1].
@@
The cloud mask provided by S2 products is prone to errors. Correspondingly, it served for data sampling only. Following [3], we used the SEnSeI v2 model [2] to provide accurate cloud masks for SSL4EO-S12 v1.1.
@@
Furthermore, we provide S2 RGB data based on L2A products. For reasons of data compression, reflectance value ranges of 0–2000 get mapped to 0–255 (single-byte integers). However, this approach may lead to saturation in bright regions (e.g., clouds, snow, and sand). To address this issue, we first remove the offset from the data. Next, we determine the 2% and 98% quantiles over the RGB channels within each image and apply a factor of 0.5 to values outside these bounds to reduce extreme pixel intensities. We then map these adjusted values to an 8-bit (0–255) range, using the 0.2% and 99.8% quantiles as the lower and upper limits, respectively. If the computed upper limit falls below 2000, we clamp it to 2000. Finally, if the median of the RGB values in an image is below 1000, we set the lower limit to 0 in order to avoid further darkening already dim images. The exact transformation is privided in Listing LABEL:lst:rgb.
Listing 1: Pseudocode for the S2-to-RGB transformation.
@@
This pipeline ensures a more visually balanced and artifact-free RGB representation of the S2 data, as visible in Figure 1.
For efficient data access and storage, we shuffled and stacked the samples into Zarr files (ZipStore, Zarr version 2Zarr documentation: https://zarr-specs.readthedocs.io/en/latest/v2/v2.0.html) in chunks of 64 samples and four timestamps each. This approach reduces the total file count, which is often important for storing data on an HPC cluster or HuggingFace. Using Zarr Zip files, the data remains compatible with the common tool xarray and enables lazy data loading via dask. The data is chunked by timestamp to enable efficient loading of individual time steps if no time series is needed. The S2 bands are stored as 16-bit (2 bytes) integers, and the S1 data is stored as 16-bit floats. Combined with Zarr’s built-in compression, these measures significantly reduce storage overhead.
@@ 4 Dataset
S1 GRD float16 -50 – +1 dB 2
S2 L1C int16 1k – 11k DN 13
S2 L2A int16 1k – 11k DN 12
S2 RGB uint8 0 – 255 3
Like v1, the dataset covers all four seasons, including snowy imagery. Sampling within a 50 km radius of the world’s 10,000 most populated cities ensures near-global coverage while focusing on urban areas but also includes ocean patches and rural areas. This diversity makes EO FMs pre-trained on SSL4EO-S12 v1.1 suitable for numerous downstream tasks, such as land use and change detection, segmentation of floods, or object detection for urban planning. The dataset also supports the training of multi-temporal models as well as multimodal fusion techniques with S1 and S2 spatially aligned. Table 1 includes further details about the modalities. Providing S2 L1C and S2 L2A ensures that pre-trained models can be used for many downstream tasks regardless of top-of-atmosphere (L1C) or bottom-of-atmosphere information (L2A).
@@
Relevant metadata is stored with the image data in the Zarr files as additional data variables. The metadata includes center coordinates in latitude and longitude (EPSG:4326) and coordinates of the corresponding UTM zones, the file identifiers, and timestamps as defined by ESA. The SEnSeI v2 cloud mask includes classes: land, water, snow, thin cloud, thick cloud, cloud shadow, and no data. The shared storage in one Zarr file simplifies data management and allows users to efficiently reference image location and timestamps.
SSL4EO-S12 v1.1 is released under the CC-BY-4.0 license, which permits broad adoption and reuse in both academic and industrial environments.

### v2

@@ Abstract
This technical report presents SSL4EO-S12 v1.1, a multimodal, multitemporal Earth Observation dataset designed for pretraining large-scale foundation models. Building on the success of SSL4EO-S12, this extension addresses the previous challenges of data misalignment and a limited data structure for low-barrier, analysis-ready EO processing. SSL4EO-S12 v1.1 covers the world’s 10,000 largest cities and its surroundings within a 50 km radius across four seasons, resulting in a diverse collection of nearly one million patches. SSL4EO-S12 v1.1 packages the data in Zarr file format for cloud-efficient loading and representation of meta-information such as including cloud masks and geolocation. Released under the CC-BY-4.0 license, SSL4EO-S12 v1.1 facilitates open research and provides a robust foundation for future advancements in self-supervised learning and geospatial analysis. The dataset is available online through https://datapub.fz-juelich.de/ssl4eo-s12, and we provided additional resources at https://github.com/DLR-MF-DAS/SSL4EO-S12-v1.1.
footnotetext: Co-funded by the European Union (Horizon Europe, Embed2Scale, 101131841). We also thank Yi Wang, main architect of the SSL4EO-S12 dataset for valuable discussions.
@@ 1 Introduction
Figure 1: Example patches from SSL4EO-S12 v1.1 showing four columns with timestamps of Sentinel-2 (S-2) L2A (left) and Sentinel-1 (S-1) GRD (right) products. S-1 GRD is visualized using a VH-$V$-VH pseudo coloring. $V$ is scaled from -30 – 5 db while VH is scaled from -40 – 0 db to account for their value ranges.
@@
Despite its success, the original SSL4EO-S12 dataset presented two main challenges: (1) misalignments between Sentinel-1 (S-1) and Sentinel-2 (S-2) modalities and (2) limited availability of analytics-ready data (ARD). To address these issues, we present SSL4EO-S12 v1.1. We ensure proper alignment by downloading larger areas and reprojecting S-1 GRD data to match S-2 L2A coordinates. Furthermore, we introduce stricter filtering and provide the dataset in an ARD format suitable for direct pretraining of EO FMs. Figure 1 illustrates S-2 L2A and S-2 GRD time series samples from SSL4EO-S12 v.1.1.
This technical report is organized as follows: We first detail the data processing steps necessary to construct SSL4EO-S12 v1.1, including alignment, reprojection, and filtering. We then present a overview of the final dataset and discuss its applicability for EO FM training pipelines.
@@ 2 Related Work
SSL4EO-S12 addressed many of these limitations by including multiple modalities (e.g., S-1 and S-2) and covering multiple seasons, thereby supporting a broader range of downstream tasks [10]. Several state-of-the-art models have leveraged SSL4EO-S12 to achieve significant improvements in accuracy and generalizability [10, 9, 4]. For instance, contrastive learning approaches and transformer-based architectures have demonstrated notable performance gains when pre-trained on SSL4EO-S12.
@@ 3 Data Processing
All data for SSL4EO-S12 v1.1 were acquired from Google Earth Engine [5]. In SSL4EO-S12, all samples were projected onto EPSG:4326, which can distort distances, especially near the poles. We retain the native UTM zones in v1.1 to preserve spatial accuracy and minimize unnecessary reprojections. For each location in v1, four timestamps from different seasons were selected from 2019 to 2021 with cloud coverage below 10% [10]. We selected the timestamp closest to the S-2 timestamps in v1. To ensure correct spatial alignment, we first downloaded a larger area of interest for both S-2 and S-1 to buffer spatial mismatch on reprojection of these modalities when paired with other datasets. Next, we checked that all S-2 timestamps for a location shared the same UTM zone. If timestamps spanned multiple zones—common at UTM boundaries–we picked the majority CRS for that location, cropped a reference timestamp to a common size of 264x264 pixels, and reprojected the remaining timestamps accordingly. Subsequently, we projected S-1 patches onto the curated 264x264-pixel S-2 patches. Following v1 [10], we fixed the patch sizes to 264 pixels as it is divisible by six, accommodating S-2 pixel resolutions of 10 m, 20 m, and 60 m. All bands were upsampled to 10 m using nearest interpolation.
@@
Next, we dropped any location with more than 1% of missing values (NaN) for any timestamp or channel. The remaining NaNs were filled in using nearest neighbor interpolation. We split the dataset into training and validation sets, following a 99%-to-1% partition defined by the TerraMesh datasetTerraMesh is a larger pretraining dataset to be released in Q2 2025. to avoid spatial overlap that may leak information while training models. We add a numerical offset of +1000 to the data to facilitate the latest processing standard defined by ESA [1].
@@
The cloud mask provided by S-2 products is prone to errors. Correspondingly, it served for data sampling only. Following [3], we used the SEnSeI v2 model [2] to provide accurate cloud masks for SSL4EO-S12 v1.1.
@@
Furthermore, we provide S-2 RGB data based on L2A products. For reasons of data compression, reflectance value ranges of 0–2000 get mapped to 0–255 (single-byte integers). However, this approach may lead to saturation in bright regions (e.g., clouds, snow, and sand). To address this issue, we first remove the offset from the data. Next, we determine the 2% and 98% quantiles over the RGB channels within each image and apply a factor of 0.5 to values outside these bounds to reduce extreme pixel intensities. We then map these adjusted values to an 8-bit (0–255) range, using the 0.2% and 99.8% quantiles as the lower and upper limits, respectively. If the computed upper limit falls below 2000, we clamp it to 2000. Finally, if the median of the RGB values in an image is below 1000, we set the lower limit to 0 in order to avoid further darkening already dim images. The exact transformation is provided in Listing LABEL:lst:rgb.
Listing 1: Pseudocode for the S-2-to-RGB transformation.
@@
This pipeline ensures a more visually balanced and artifact-free RGB representation of the S-2 data, as visible in Figure 1.
For efficient data access and storage, we shuffled and stacked the samples into Zarr files (ZipStore, Zarr version 2Zarr documentation: https://zarr-specs.readthedocs.io/en/latest/v2/v2.0.html) in chunks of 64 samples and four timestamps each. This approach reduces the total file count, which is often important for storing data on an HPC cluster or HuggingFace. Using Zarr Zip files, the data remains compatible with the common tool xarray and enables lazy data loading via dask. The data is chunked by timestamp to enable efficient loading of individual time steps if no time series is needed. The S-2 bands are stored as 16-bit (2 bytes) integers, and the S-1 data is stored as 16-bit floats. Combined with Zarr’s built-in compression, these measures significantly reduce storage overhead.
@@ 4 Dataset
S-1 GRD float16 -50 – +1 dB 2
S-2 L1C int16 1k – 11k DN 13
S-2 L2A int16 1k – 11k DN 12
S-2 RGB uint8 0 – 255 3
Like v1, the dataset covers all four seasons, including snowy imagery. Sampling within a 50 km radius of the world’s 10,000 most populated cities ensures near-global coverage while focusing on urban areas but also includes ocean patches and rural areas. This diversity makes EO FMs pre-trained on SSL4EO-S12 v1.1 suitable for numerous downstream tasks, such as land use and change detection, segmentation of floods, or object detection for urban planning. The dataset also supports the training of multi-temporal models as well as multimodal fusion techniques with S-1 and S-2 spatially aligned. Table 1 includes further details about the modalities. Providing S-2 L1C and S-2 L2A ensures that pre-trained models can be used for many downstream tasks regardless of top-of-atmosphere (L1C) or bottom-of-atmosphere information (L2A).
@@
Relevant metadata is stored with the image data in the Zarr files as additional data variables. The metadata includes center coordinates in latitude and longitude (EPSG:4326) and pixel coordinates in the corresponding UTM zones, the file identifiers, as well as timestamps as defined by ESA. The SEnSeI v2 cloud mask includes classes: land, water, snow, thin cloud, thick cloud, cloud shadow, and no data. The shared storage in one Zarr file simplifies data management and allows users to efficiently reference image location and timestamps.
SSL4EO-S12 v1.1 is released under the CC-BY-4.0 license, which permits broad adoption and reuse in both academic and industrial environments. SSL4EO-S12 v1.1 adds to the family of SSL4EO-$\ast$, open-source datasets to serve the Earth observation data science community.